OpenStudy (anonymous):

Determine if the graphs of the pair lines are parallel X+7=y y-x=-3

OpenStudy (aroub):

no

OpenStudy (saifoo.khan):

Yes they are!

OpenStudy (anonymous):

ok so here's what you do: 1) move the Y's to one side and the constants and X's to the other. 2) to check for "parallel", look at the number in front of X. if they are the same, (including negatives), then they are parallel!

OpenStudy (anonymous):

x+ 7 = y Y = x + 7 y-x = -3 Y = x -3 so now we have the two graphs: Y = x + 7 and Y = x -3 To compare each and decide whether they are parallel we look at the SLOPE of each graph. This is the variable in front of X in the equations. Since they are both ONE that means they have the same SLOPE and therefore are PARALLEL lines.

OpenStudy (aroub):

yes sorry they are y=7+x,y=x+3 same slopes so they are parallel